# Kinetic coupling and dark matter on the lattice K. Farakosa, V.Iconomoua and G. Koutsoumbasa

### v1

@@ 5 Conclusions
(5) The system may move from the broken to the symmetric phase, when is changed. According to figure 10, there is a phase transition from the broken to the symmetric phase in model B, as becomes bigger than 1.75. This explains the qualitatively different behaviour for in figures 8 and 9. Similarly, according to figure 12, there is a phase transition from the broken to the symmetric phase in model C, as becomes bigger than 1.63. This explains the qualitatively different behaviour for in figure 13. The critical values of move to smaller values (for instance 0.67 or 0.18, as depicted in figure 12) if is chosen to be smaller, i.e. for stronger gauge coupling.
@@
Future work may include simulation of a model based on to provide the possibility to study the main physically different and realistic couplings of the dark to either the hypercharge or the electromagnetic

### v2

@@ 5 Conclusions
(5) The system may move from the broken to the symmetric phase, when is changed. According to figure 10, there is a phase transition from the broken to the symmetric phase in model B, as becomes bigger than 1.75. This explains the qualitatively different behaviour for in figures 8 and 9. Similarly, according to figure 12, there is a phase transition from the broken to the symmetric phase in model C, as becomes bigger than 1.63. This explains the qualitatively different behaviour for in figure 13. The critical values of move to smaller values (for instance 0.67 or 0.18, as depicted in figure 12) if is chosen to be smaller, i.e. for stronger gauge coupling. In an attempt to relate lattice parameters with continuum quantities we also write down the effective potential in the continuum at one loop, where the gauge coupling should be replaced by with a quantity independent from This form describes the effective coupling in the presence of kinetic coupling, as explained in the paper; it is expected to provide valuable non-perturbative information. The result is that the characteristic of two minima persists, as one increases however the separation of the two minima becomes very small. This means that the phase transition towards the symmetric phase is missed by the one-loop effective potential, but it appears that the system will possibly move to the symmetric phase near if higher loop corrections are included; we are working on this.
@@
Future work may include:
(a) Numerical computation for the functions for the various couplings via a non-perturbative calculation.
111We thank the referee for raising this issue.
To this end one might compute the dependence of our results on the lattice size, to get an estimate of the dependence on the lattice spacing and the approach to the thermodynamic limit. This calculation would allow for a connection with physical quantities and we postpone it for a separate publication.
(b) Simulation of a model based on to provide the possibility to study the main physically different and realistic couplings of the dark to either the hypercharge or the electromagnetic